# Hydrodynamics of the electronic Fermi liquid: a pedagogical overview

### v1

@@ III.3 Electron-Phonon Hydrodynamics
So far we have described electron-phonon scattering merely as a source of momentum relaxation, such that phonons take momentum from the electron system and dissipate it elsewhere. As such, we have considered such scattering as detrimental to the formation of a hydrodynamic electron fluid. However, if the phonons relax momentum slowly, they can also transfer momentum back into the electron system. As such, the combined electron-phonon system could be momentum-conserving and be in a hydrodynamic regime, even though the electron component alone may not be hydrodynamic. Analogous to the electron-hole fluid, the electron-phonon fluid is comprised of two quasiparticles, electrons and phonons. Just as before, one can use the same Boltzmann approach to derive continuity equations and constitutive relations. As the mathematical formalism is much the same, we only briefly comment on this regime and leave detailed description of the results to Refs. 64, 214, 9.
The most dramatic qualitative effect of electron-phonon hydrodynamics is phonon drag – if the electron-phonon fluid is strongly coupled, the electrons and phonons will tend to move together. Thus, upon applying an electric field, the charged electrons experience a driving force and drag the neutral phonons along. On the other hand, a temperature gradient produces a driving force for both electrons and phonons. This asymmetry leads to violations of the WF law, as well as an enhancement of thermoelectric effects. However, as previously mentioned for the clean Fermi liquid, observations of these effects must be done with care to exclude other possible pathways to the same effect. Experimentally, electron-phonon hydrodynamics has been proposed for Sb [215], WTe2 [48], WP2 [47, 46], ZrTe5 [216], PtSn4 [217, 218], and NbGe2 [219], though detailed confirmations between theory and experiment have yet to be done.
@@ IV Conclusion and Outlook
The observation of the hydrodynamic regime and the treatment of electron-electron interactions have led to a reevaluation of traditional transport lore. The transport properties of a clean, hydrodynamic Fermi liquid are qualitatively different in many regards from the usual momentum-relaxing metallic Fermi liquid. For instance, even the basic concept of a local resistivity fails in the hydrodynamic material and the “fundamental” Landauer bound on conductance is less fundamental than originally thought. Such phenomena have motivated not only an interest in novel materials, but also a resurgence of interest in long-studied 2DEGs with a more careful eye towards hydrodynamic effects [26, 38, 39, 40, 41, 42, 43, 44, 45].
Throughout this review, we have characterized the hydrodynamic regime as “strongly-interacting.” However, we have mostly contented ourselves with the regime where interactions are strong relative to momentum-relaxing scattering processes but still sufficiently weak that a Boltzmann description still applies, i.e. the interactions can still be treated perturbatively. To draw an analogy to classical fluids, these are “gas-like” systems. However, there are also many “liquid-like” systems where interactions are non-perturbatively strong, e.g. systems where quasiparticles are short-lived. Even in the classical case, the behavior of liquids can be very different from those of gases. For instance, viscosity tends to decrease with temperature in a classical liquid while viscosity increases for a classical gas. This observation has led to classical conjectures on the possibility of a lower bound on viscosity [220, 221] and a quantum counterpart originating from the AdS/CFT literature [222]. One could also hope that hydrodynamics may provide a new window into long-standing transport mysteries. The strongly correlated strange metal phase is one such example, which has a low resistivity $\rho\sim T$ that mysteriously scales linearly with $T$ [223]. Several works have suggested links to hydrodynamic behavior [224, 225, 226, 227, 123], though its true origin is still unclear. The power of the hydrodynamic framework is to unify, in a non-perturbative way, the language with which we discuss transport behavior in a wide array of systems. In so doing, we not only uncover novel transport phenomena but hope that this framework may shed some light on long-standing mysteries in strongly correlated transport.
@@ V Acknowledgments
We thank Patrick Ledwith for helpful comments regarding this review. We thank Yihang Zeng, Cory Dean, and Felicia Setiono for graciously providing figures. B.S. was supported by the NSF under Grant No. DMR-2045742.

### v2

@@ III.3 Electron-Phonon Hydrodynamics
So far we have described electron-phonon scattering merely as a source of momentum relaxation, such that phonons take momentum from the electron system and dissipate it elsewhere. As such, we have considered such scattering as detrimental to the formation of a hydrodynamic electron fluid. However, if the phonons relax momentum slowly, they can also transfer momentum back into the electron system. As such, the combined electron-phonon system could be momentum-conserving and be in a hydrodynamic regime, even though the electron component alone may not be hydrodynamic. Analogous to the electron-hole fluid, the electron-phonon fluid is comprised of two quasiparticles, electrons and phonons. Just as before, one can use the same Boltzmann approach to derive continuity equations and constitutive relations. As the mathematical formalism is much the same, we only briefly comment on this regime and leave detailed description of the results to Refs. 85, 237, 9.
The most dramatic qualitative effect of electron-phonon hydrodynamics is phonon drag – if the electron-phonon fluid is strongly coupled, the electrons and phonons will tend to move together. Thus, upon applying an electric field, the charged electrons experience a driving force and drag the neutral phonons along. On the other hand, a temperature gradient produces a driving force for both electrons and phonons. This asymmetry leads to violations of the WF law, as well as an enhancement of thermoelectric effects. (Hydrodynamic phonons interacting with electrons have also been long suggested to produce interesting forms of second sound [238, 239].) However, as previously mentioned for the clean Fermi liquid, observations of these effects must be done with care to exclude other possible pathways to the same effect. Experimentally, electron-phonon hydrodynamics has been proposed for Sb [240], WTe2 [64], WP2 [63, 62], ZrTe5 [241], PtSn4 [242, 243], and NbGe2 [244], though detailed confirmations between theory and experiment have yet to be done.
@@ IV Conclusion and Outlook
The observation of the hydrodynamic regime and the treatment of electron-electron interactions have led to a reevaluation of traditional transport lore. The transport properties of a clean, hydrodynamic Fermi liquid are qualitatively different in many regards from the usual momentum-relaxing metallic Fermi liquid. For instance, even the basic concept of a local resistivity fails in the hydrodynamic material and the “fundamental” Landauer bound on conductance is less fundamental than originally thought. Such phenomena have motivated not only an interest in novel materials, but also a resurgence of interest in long-studied 2DEGs with a more careful eye towards hydrodynamic effects [42, 54, 55, 56, 57, 58, 59, 60, 61]. Though many of the effects described have a classical fluids analogue, we have noted some non-trivial extensions of hydrodynamic behavior in electronic systems in Sec. III. Furthermore, one can also extend the hydrodynamic or kinetic theory approach to include quantum or topological effects, such as Berry curvature [245, 246].
Throughout this review, we have characterized the hydrodynamic regime as “strongly-interacting.” However, we have mostly contented ourselves with the regime where interactions are strong relative to momentum-relaxing scattering processes but still sufficiently weak that a Boltzmann description still applies, i.e. the interactions can still be treated perturbatively. To draw an analogy to classical fluids, these are “gas-like” systems. However, there are also many “liquid-like” systems where interactions are non-perturbatively strong, e.g. systems where quasiparticles are short-lived and the Boltzmann approach fails. Even in the classical case, the behavior of liquids can be very different from those of gases. For instance, viscosity tends to decrease with temperature in a classical liquid while viscosity increases for a classical gas. This observation has led to classical conjectures on the possibility of a lower bound on viscosity [247, 248] and a quantum counterpart originating from the AdS/CFT literature [249]. In general, strongly-correlated (“liquid-like”) systems are difficult and nigh-intractable to study from a first-principles approach.
The surprising power of the hydrodynamic framework is that it works equally well for classical liquids as it does for classical gases, despite the inability to derive the equations, e.g. from a Boltzmann framework. One could also hope that the hydrodynamic framework may provide a new window into long-standing transport mysteries. The strongly correlated strange metal phase is one such example, which has a low resistivity $\rho\sim T$ that mysteriously scales linearly with $T$ [250]. Several works have suggested links to hydrodynamic behavior [251, 252, 253, 254, 144], though its true origin is still unclear. The power of the hydrodynamic framework is to unify, in a non-perturbative way, the language with which we discuss transport behavior in a wide array of systems. In so doing, we not only uncover novel transport phenomena but hope that this framework may shed some light on long-standing mysteries in strongly correlated transport.
@@ V Acknowledgments
We thank Patrick Ledwith, Sankar Das Sarma, and Boris Shklovskii for helpful comments regarding this review. We thank Yihang Zeng, Cory Dean, and Felicia Setiono for graciously providing figures. B.S. was supported by the NSF under Grant No. DMR-2045742.